The n-th harmonic number is defined by



Hn=∑n, i=1 (1/i) =1+1/2+1/3+…+1/n.



Show that



Hn=∫(1>0) 1−x^n/1−x dx.

1
Expert's answer
2021-11-15T16:44:22-0500

"x=1-u"


"\int^1_0\frac{1-x^n}{1-x}dx=-\int^0_1\frac{1-(1-u)^n}{u}du=\int^1_0\frac{1-(1-u)^n}{u}du="


"=\int^1_0[\displaystyle{\sum_{k=1}^n(-1)^{k-1}\begin{pmatrix}
 n \\
 k 
\end{pmatrix}u^{k-1}}]du=\displaystyle{\sum_{k=1}^n(-1)^{k-1}}\begin{pmatrix}
 n \\
 k 
\end{pmatrix}\int^1_0u^{k-1}du="


"=\displaystyle{\sum_{k=1}^n(-1)^{k-1}}\frac{1}{k}\begin{pmatrix}
 n \\
 k 
\end{pmatrix}"